y^3/4
_____
y^1/4

Simplify...

I keep forgetting the first step, any idea?

Answer by stanbon(75887)

y^3/4
_____
y^1/4
------------
When dividing, subtract the exponents:
= y^((3/4)-(1/4))
= y^(1/2)
= sqrt(y)
==================
Cheers,
Stan H.